A 12.62 g sample of a compound containing only C,H and O produces 27.73 g of CO2 and 9.09 g H2O in a combustion analysis. Its molecular mass is found to be 300 g. Determine its a) empirical formula b) molecular formula


1
Expert's answer
2020-11-17T10:39:56-0500

Molar mass of CO2 = 44.01 g/mol

Molar mass of H2O = 18.01 g/mol

mass of C in CO2 = "(m \u3016CO\u3017_2 \u00d7 M of C)\/(M \u3016CO\u3017_2 )= (27.73 g \u00d712.01 g\/mol)\/(44.01 g\/mol)"

= 7.57 g C

mass of H in H2O = "( m H_2 O \u00d72 \u00d7M of H)\/(M H_2 O) = (9.09 g \u00d72\u00d71.01)\/(18.01 g\/mol)"    = 1.02 g H

mass of O = 12.62g - 7.57g - 1.02g = 4.03 g O

7.57g C×(1mol C)/12.01g C) = 0.63 mol C

1.02g H×(1mol H)/(1.01g H) = 1.0 mol H

4.03g O×(1mol O)/(16.00g O) = 0.2525 mol O

Since atoms combine in the same ratio that moles do, we divide all of the numbers of moles by the smallest number to put everything into lowest terms:

C0.63 H1.0 O0.2525 → C0.63/0.2525 H1.0/0.2525 O0.2525/0.2525 → C2.5 H4.0 O1.0

If the mole ratio is not all whole numbers, we multiply through by the smallest integer which will turn all of the numbers into integers. These numbers are the subscripts of the elements in the empirical formula.

(C2.5 H4.0 O1.0)×2 = C5 H8 O2 (empirical formula)

If we know the molar mass of the compound, we can obtain the molecular formula by dividing the weight of the empirical formula into the molar mass; this will determine the number of empirical formula units in the molecule.

(C5 H8 O2 )× 3 = C15H24O6

Answer: empirical formula C5H8O2; molecular formula C15H24O6